# labelinglabel

### v1

@@ 2 The Barrow universe’s essence is nothing but Einstein quintessence
[…]
@@
Since we are talking about a collapsing universe, let us set ${\dot{a}}<0$. The general solution of (7) for a closed unverse with $w=-1/3$ has the form:
@@
which explains the term “Einstein quintessence” chosen for this model (see [14]). Hence, a zero universe (3) is indeed a special case of EQ (11) with $a_{s}=0$! But is it the only model where it might naturally arise? Can we not gain a zero universe (3) by some other means besides the reduction from EQ? To answer this question let us first reiterate that the open universe case (when $k=-1$) must be discarded right away, owing to the requirement $\rho\geq 0$. This leaves us the flat universes with $k=0$. For them both ${\dot{a}}\to 0$ and $a\to 0$ only when $-1<w<-1/3$ (the phantom case $w<-1$ deserves a separate discussion, and will be granted one in Sec. 6). Integrating (7) leads to:
@@ 3 Barrow Universe and The Multiverse of Sanity
[…]
@@
. The fields of matter in the rest of the universes with $n>2$ satisfy the standard energy conditions employed in the theorems about the singularities [18]. In particular, we assume that $\rho_{n}$ has but one special point that occurs only when $a_{n}\to 0$.   [^9] This implies that by using this assumption we are automatically dismissing the possible singularities arising for either a finite non-zero or infinite values of scale factor, such as the Big Rip singularity [19], [20],[21],[22],[23],[24], the Big Freeze singularity [25], [26],[27], the singularities localized on a brane [28], the Sudden Future singularity [29], [30],[31], [24], [32], [33], the Big Boost singularity [34], and the Big Break singularity [35], [36]. For a useful classification (N.O.T.) of those singularities we refer the reader to the article [37]. Actually, some of them, such as a Sudden singularity (with a finite value of the first derivative os scale factor) might still be included into our consideration. However, for the purposes of this article we are deliberately excluding almost anything that can be considered “exotic”, restricting ourselves solely to the phenomena produced by the quantum potential (14). For that end, we are also abstaining from the sudden-like singularity from the [38] and from the unusual types of singularities that manifests themselves as finite jump discontinuities of cosmological parameters [39]. The only exception are the Big Rip singularities which will be dealt with in Sec. 6.
@@ 4 The Conspicuous Absence of Singularities
Before we begin formulating the main Theorem that rules out singularities in the proposed model, it will be a good idea to get a bit more acquainted with the quantum potential (14). It depends on $N$ scale factors $a_{n}$, whose appearances follow a very simple pattern: every $a_{n}$ with indices $12<n<N-$ 1 is present in just three terms:
$\begin{split}U(a_{1},...,{\bf a_{n}},...a_{{}_{N}})=...&+\left(\frac{1}{{\bf a_{n}}-a_{n-1}}-\frac{1}{a_{n-1}-a_{n-2}}\right)^{2}\\ &+\left(\frac{1}{a_{n+1}-{\bf a_{n}}}-\frac{1}{{\bf a_{n}}-a_{n-1}}\right)^{2}+\\ &+\left(\frac{1}{a_{n+2}-a_{n+1}}-\frac{1}{a_{n+1}-{\bf a_{n}}}\right)^{2}+....\end{split}$ (24)
@@
[…]
For the sake of simplicity let $t^{\prime}=0$  [^10] This can be done without any loss of generality thanks to the invariance of (16) and (17) to the time translations and time inversions. From the point of view of functions $\{a_{n}\}_{n=2}^{N}$ there are two ways to satisfy (26) and (27) at $t=0$: (i) as $t\to 0$ we should have $a_{k}/a_{k+1}\to 0$ for $k=2,..,n-1$ or (ii) $a_{k}/a_{k+1}\to{\rm const}$. In the former case in the close vicinity of $t=0$ the functions $a_{n}$ behave as a set of power functions:
@@
where $\mu_{k}={\rm const}$ and the exponents $m_{k}\in(0,1)$ in order to satisfy the Conditions 1 and 4. Case (ii) is essentially a variation of a “master-factor hypothesis” from [8], localized in a sufficiently small vicinity of $t=0$, and it implies the existence of such function $a(t)$ and such dimensionless constants $\mu_{k}$, $k=1,...,\,n$, that
@@
where $\epsilon^{2}$ and $t_{0}$ are two real-valued constants of integration. (32) unequivocally proves that $a_{2}(t)$ cannot turn to zero for any value of $t$, directly contradicting our initial assumption and thus proving our theorem for the Case (i)  [^11] It is also possible to prove our theorem by utilizing (28) directly; the direct computation $10<m_{2}<$ 1 would yield the analogue of (31) where the l.h.s. is always negative whereas the r.h.s. is strictly positive, thus producing the thought after discrepancy.
@@ 5 The Taming of the Few: Some Exact Solutions
  • The purely quantum universe: $\rho(x)=0$. In this case (33) has a real-valued solution only for the open model, $k=-1$. The solution itself has the form:
@@
    (ii) An open model, $k=-1$. In this case the solution has the following parametric form:
@@
    ±$\begin{split}\pm t=t_{0}+&\sqrt{\frac{2}{\mu c^{2}}}\left(\frac{1}{3}\left(x-\frac{2L_{{}_{PL}}^{2}}{\mu}\right)^{3/2}+\right.\\ &\left.+\frac{6L_{{}_{PL}}^{2}}{\mu}\left(x-\frac{2L_{{}_{PL}}^{2}}{\mu}\right)^{1/2}\right).\end{split}$ (38)
    Once again, (38) has no singularities ($x_{{\rm min}}=2L_{{}_{PL}}^{2}/\mu$), and at $02L_{{}_{PL}}=$ 0 reduces to a familiar solution: $x\propto t^{2/3}$.
@@
    Interestingly, this implies that, quite unlike the dust dominated universe (36), the period of oscillation now actually depend upon the quantum effects (albeit in an admittedly almost imperceptible manner). The solution (42) with $k=-1$ describes an open universe emerging from an initial singularity at $t=2\sqrt{\delta}/c^{2}$, and (43) deals with a flat universe, which is also born from a singularity at $t=0$.
    (ii) $\delta=0$. This case as well the next one can only be realized in the presence of a quantum potential. The equation (33) reduces to ${\dot{x}}^{2}+kc^{2}=0$ which has real-valued solutions either when $k=-1$ ($x(t)=\pm c(t-t_{0})$) or $k=0$ ($x={\rm const}$).
@@
    and the universe must be necessarily open ($k=-1$).
@@ 5.2 Three is a company: $N=3$ and the Hamilton-Jacobi equation
Now, what kind of fields of matter should fill the third universe? In this subsection we will consider three cases: (I) there are no fields of matter ($\rho_{3}=p_{3}=0$) and the third universe is, like the second one, governed purely by a quantum potential; (II) a universe filled with matter whose parameter of state is $w=-1/3$ (and so $\rho_{3}=\sigma^{2}/a_{3}^{2}$); (III) a hot universe filled by radiation with $w=1/3$ (here $\rho_{3}=R^{2}/a_{3}^{4}$). The reason we chose them specifically is because in these three cases the equations turns out to be completely integrable and leading to solutions with a strikingly similar behaviour. It is crucial to keep in mind that unlike the previously discussed cases, when $N>2$ the equations (16), (17) generally cease to be completely integrable, so the importance of those rare cases with known exact solutions cannot be overestimated. In fact, in the remainder of this section we will discuss both the solutions and the method of obtaining them.
@@
[…]
@@
The behaviour of the solution (56) explicitly depends upon the choice of $v(\phi)$ and will be different in cases (I), (II) and (III). On the other hand, we can still surmise asymptotic behaviour of the universes #$\#2$ and #$\#3$ without actually taking the integrals in (56). Indeed, using (47) we can calculate the Hubble parameter:
@@
and $v(\phi)$ is given by (50). Note, that except for the pathologically small values of constant $R\to 0$, the function $-v(\phi)$ remains strictly positive in between the stop points $\phi_{1}$ and $\phi_{2}$ that satisfy the condition $40<\phi_{1}<\phi_{2}<\pi/$ 4, thus making sure the solutions (60) remain real valued.

### v3

# Quantum Cosmology Without Singularities: A New Approach

@@ Abstract
Preface
A reader might wonder: why start a scientific article with a Preface – a mode of introduction seldomly used outside of anthologies, monographes and second editions? If the authors have something important to spill, why not pour it all into the Introduction, as is common and proper?.. The answer is rather simple. Just as this paper was receiving its final polish, we, the authors, have reached a chilling realization: in our quest of explaining all the methods and reasonings in as many details as possible (so that even an inattentive reader would not struggle too much following the article’s logic), we actually overshoot by not providing a potential reader with a brief, straightforward explanation as to what this article actually is and isn’t about. This realization was very troubling because, as everyone knows, a modern reader is all too eager to dump an unread article into an unreliable LLM chatbot, only to receive a rather questionable summary of a dubious value, and then regard the original article as “banal and ”. So, we had no choice but to write this Preface!
[…]
The key results of the paper can be summed up as follows:
  1. 1.
    The quantization of cosmological solutions via the Hall, Deckert and Wiseman approach appears to be better suited for the expanding Friedman universes than the standard quantization methods, owing to the presence of cosmological horizons. The method becomes especially relevant for de Sitter (dS) universes, where the horizon is a key cosmological features – in fact, for dS the MIW approach might be the only logical one (see Sec. 1.2).
  2. 2.
    While an early Friedman universe is effectively “smoothed over” by inflation, on the later stages it might experience a direct manifestation of quantum effects on a scale of a cosmological horizon. It is important to note, that we are not referring here to a well known phenomena of quantum fluctuations being “blown up” by inflation and subsequently evolving into contemporary galaxies. What we are claiming instead is that the quantum effects can potentially influence a dynamical expansion of the universe, and might therefore be observable not just on an atomic scale, but on a cosmological scale (about $10^{28}$ cm) as well (see also [2]). In order to avoid a possible confusion, let us stress here that the quantum phenomena do not manifest themselves on the intermediate scales (between the micro and the mega), where a quantum decoherence reigns supreme.
  3. 3.
    At the core of our approach lies a model with a set (finite or countable) $\mathcal{D}$ of universes interacting via the Hall-Deckert-Wiseman quantum potential (see Appendix A for its explanation, derivation, and relationship to a famous de Broglie-Bohm potential). As we shall soon see, by (i) adding to this set a Barrows’s zero universe (a seemingly ill-defined object of zero scale factor [3]), complimented with (ii) an additional empty universe with no fields of matter, and then (iii) by imposing a few simple physically sensible assumptions (see Sec. 3, conditions 1-3), a most unusual Theorem can be proven – that the Hall-Deckert-Wiseman quantization automatically eliminates the singularities with zero values of a scale factor (see Sec. 4, Theorem 1). For instance, a collapsing universe on an approach to a Big Crunch singularity experiences an effective rebound and undergos a new phase of expansion; for this no additional assumptions or fields of matter are required.
  4. 4.
    Additionally, a quantization under the aforementioned three conditions automatically gets rid of any Big Rip singularities (BR). In other words, any universe from $\mathcal{D}$ containing phantom fields with arbitrary equation of state (as long as w<1w<-1) are effectively shielded from experiencing BR. This remarkable property is a purely kinematic phenomena and is a Corollary of the aforementioned Theorem (see Sec. 6, Theorem 2).
  5. 5.
    Although on a first glance a Barrow universe with zero scale factor [3] seems to be ill-posed, it is indeed an exact solution of Einstein equations. Furthermore, this solutions is merely a special case of the Einstein quintessence model, which describes a closed stationary Einstein universe (see Sec. 2) – and is itself a special case of a Tipler’s Omega point solution. The fact that a zero universe is a full-fledged (albeit exotic) exact solution means that it has to be accounted for during the quantization procedure, be it via the functional integral method or by the Hall-Deckert-Wiseman approach. To put it in other words, an addition of the “Barrow universe” into a list of interacting universes is neither artificial nor a whim, but a natural and even essential step for a quantization of gravity.
Again, we emphasize that these five statements are not hypothetical – they are the results derived and (in case of the two theorems) proven in the current paper. However, we might add one more statement to this list, admittedly a more speculative one, but which nevertheless appears to be a logical conclusion of the Hall-Deckert-Wiseman formalism applied to an early universe:
  1. 6.
    The MIU method appears to provide interesting solutions for two fundamental problems of inflationary theory: for the decoherence conundrum [4] and for the mystery of extremely low initial entropy, indispensable for jump-starting the process of inflation [5], [6].
Again, we must warn the reader that we don’t have the same rigorous proof for the last statement as we do for the previous five. However, as we discuss in Sec. 7, there are some evidences in favour of this statement.
Before we wrap the Preface up, in the spirit of completeness we would like to adumbrate the limitations of both the method and the results of this paper.
  • The paper is dedicated solely to singularities with zero values of scale factor, with an exception of Sec. 6, which throws the Big Rip singularities into the mix. In order to achieve this we impose strict energy conditions upon the types of matter that fill the interacting universes. Thus, we are automatically dismissing possible singularities arising for finite non-zero values of scale factor, such as the Big Freeze singularity [7], [8],[9], the singularities localized on a brane [10], the Sudden Future singularity [11], [12],[13], [14], [15], the Big Boost singularity [16], and the Big Break singularity [17], [18] [^5] For the readers interested in classification of these singularities, we would like to refer them to the useful article [19].. As a matter of fact, some of them – like a Sudden singularity (with a finite value of the first derivative of scale factor) – might still be included into our considerations. However, for the purposes of this article we are deliberately excluding almost anything that can be considered “exotic”, restricting ourselves solely to the phenomena produced by the quantum potential (14). For that end, we are also abstaining from the sudden-like singularity from the [20] and from the unusual types of singularities that manifests themselves as finite jump discontinuities of cosmological parameters [21]. A detailed study of such late-time singularities, which significantly differ from the traditional Big Bang/Big Crunch, has been done in [22] via the standard quantization approaches.
  • We have restricted ourselves to standard Einstein equations, without resorting to the models of modified gravity. This was done for a reason: in order to apply the MIU method of quantization to modified gravity models, the method itself must be modified, which lies far out of scope of this paper. An example of quantum cosmological models constructed via the traditional quantization recipes can be found in [23].
  • On a cosmological scale the quantum effects manifests themselves as additional forces of repulsion, contributing to an accelerated expansion of the universe. So it would be natural to wonder, whether the quantum potential might actually explain the phenomenon of dark energy in observable universe. As far as we know, the answer is likely negative: the quantum interaction between the universes does not seem to be enough to explain observable acceleration of our universe – a preliminary attempt reconciling the predictions of such a model with existing measurements on supernova apparent magnitude, Hubble parameter and baryon acoustic oscillations proved to be unsuccessful [24]. On the other hand, the quantum potential might contribute to this acceleration, which implies that the dark energy might be a compound of both cosmological constant and of the quantum potential. However, verifying this hypothesis also lies out of scope of the current paper.
And now – let’s finally get on with the show!
@@ 2 The Barrow universe’s essence is nothing but Einstein quintessence
This is, in fact, a natural extension of (1): after all, if (3) is violated, and there exists a moment of time $t_{1}$ when at least one of the higher derivatives $d^{n}a/dt^{n}$ pops out of a zero background, then there would necessarily follow a moment $t_{2}>t_{1}$, arbitrarily close to $t_{1}$, when the derivative $d^{n-1}a/dt^{n-1}$ is to become non-zero. The same fate would then befall the (n2)(n-2)-nd derivative, followed by the (n3)(n-3)-rd derivative, and so on. The resulting domino effect will necessarily culminate in a violation of the condition (1) [^9] In fact, if we assume that $a^{(n)}(t)=\epsilon(t)$ while all lower order derivatives of $a(t)$ and the scale factor itself are equal to zero at some $t=t_{1}$, then we can use the Cauchy formula for repeated integration to get the formulas: $a(t)=\frac{1}{(n-1)!}\int\limits_{t_{1}}^{t}(t-\xi)^{n-1}\cdot\epsilon(\xi)d\xi,\qquad\dot{a}(t)=\frac{1}{(n-2)!}\int\limits_{t_{1}}^{t}(t-\xi)^{n-2}\cdot\epsilon(\xi)d\xi,$ that obviously violate (3) if $\epsilon(t)\neq 0$ for $\forall t$., thus motivating and explaining our interest in the condition (3). Hence, the main goal for this section would be to demonstrate that the universes that satisfy this condition naturally arise as special solutions of the Friedmann-Lemaître-Robertson-Walker (FLRW) equations.
@@
Since we are talking about a collapsing universe, let us set ${\dot{a}}<0$. The general solution of (7) for a closed unverse with w=1/3w=-1/3 has the form:
@@
which explains the term “Einstein quintessence” chosen for this model (see [35]). Hence, a zero universe (3) is indeed a special case of EQ (11) with $a_{s}=0$.
But is it the only model where it might naturally arise? Can we not gain a zero universe (3) by some other means besides the reduction from EQ? To answer this question let us first reiterate that the open universe case (when k=1k=-1) must be discarded right away, owing to the requirement $\rho\geq 0$. This leaves us the flat universes with $k=0$. For them both ${\dot{a}}\to 0$ and $a\to 0$ only when 1<w<1/3-1<w<-1/3 (the phantom case w<1w<-1 deserves a separate discussion, and will be granted one in Sec. 6). Integrating (7) leads to:
@@ 3 Barrow Universe and The Multiverse of Sanity
where $\rho_{n}=\rho_{n}(a_{n})$ is the density of matter in the $n$-the universe (and it satisfies the usual continuity equation [2]), $p_{n}=p_{n}(a_{n})$ denotes the pressure, constants $k_{n}=-1,\,0,\,1$ are the Gaussian curvatures, and for simplicity we have introduced an additional Plank multiplier: $L_{{}_{PL}}=(8\pi\hbar G/3c^{3})^{1/2}$. Thus, in order to understand the cosmological dynamics in the resulting multiverse we have to integrate the equations (16) with the constraint (17) for the $N$ given equations of state and initial conditions.
@@
. The fields of matter in the rest of the universes with $n>2$ satisfy the standard energy conditions employed in the theorems about the singularities [39]. In particular, we assume that $\rho_{n}$ has but one special point that occurs only when $a_{n}\to 0$.   [^11] The only exception are the Big Rip singularities which will be dealt with in Sec. 6.
@@ 4 The Conspicuous Absence of Singularities
Before we begin formulating the main Theorem that rules out singularities in the proposed model, it will be a good idea to get a bit more acquainted with the quantum potential (14). It depends on $N$ scale factors $a_{n}$, whose appearances follow a very simple pattern: every $a_{n}$ with indices 2<n<N12<n<N-1 is present in just three terms:
[…]
@@
$a_{k}(t)<a_{k+1}(t),\qquad\forall\,\,{2\leq k\leq n},\qquad t\neq t^{\prime}.$ (27)
For the sake of simplicity let $t^{\prime}=0$  [^12] This can be done without any loss of generality thanks to the invariance of (16) and (17) to the time translations and time inversions. From the point of view of functions $\{a_{n}\}_{n=2}^{N}$ there are two ways to satisfy (26) and (27) at $t=0$: (i) as $t\to 0$ we should have $a_{k}/a_{k+1}\to 0$ for k=2,..,n1k=2,..,n-1 or (ii) $a_{k}/a_{k+1}\to{\rm const}$. In the former case in the close vicinity of $t=0$ the functions $a_{n}$ behave as a set of power functions:
@@
where $\mu_{k}={\rm const}$ and the exponents $m_{k}\in(0,1)$ in order to satisfy the Conditions 1 and 4. Case (ii) is essentially a variation of a “master-factor hypothesis” from [2], localized in a sufficiently small vicinity of $t=0$, and it implies the existence of such function $a(t)$ and such dimensionless constants $\mu_{k}$, k=1,,nk=1,...,\,n, that
@@
where $\epsilon^{2}$ and $t_{0}$ are two real-valued constants of integration. (32) unequivocally proves that $a_{2}(t)$ cannot turn to zero for any value of $t$, directly contradicting our initial assumption and thus proving our theorem for the Case (i)  [^13] It is also possible to prove our theorem by utilizing (28) directly; the direct computation $0<m_{2}<1$ would yield the analogue of (31) where the l.h.s. is always negative whereas the r.h.s. is strictly positive, thus producing the thought after discrepancy.
@@ 5 The Taming of the Few: Some Exact Solutions
  • The purely quantum universe: $\rho(x)=0$. In this case (33) has a real-valued solution only for the open model, k=1k=-1. The solution itself has the form:
@@
    (ii) An open model, k=1k=-1. In this case the solution has the following parametric form:
@@
    $\begin{split}\pm t=t_{0}+&\sqrt{\frac{2}{\mu c^{2}}}\left(\frac{1}{3}\left(x-\frac{2L_{{}_{PL}}^{2}}{\mu}\right)^{3/2}+\right.\\ &\left.+\frac{6L_{{}_{PL}}^{2}}{\mu}\left(x-\frac{2L_{{}_{PL}}^{2}}{\mu}\right)^{1/2}\right).\end{split}$ (38)
    Once again, (38) has no singularities ($x_{{\rm min}}=2L_{{}_{PL}}^{2}/\mu$), and at $2L_{{}_{PL}}=0$ reduces to a familiar solution: $x\propto t^{2/3}$.
@@
    Interestingly, this implies that, quite unlike the dust dominated universe (36), the period of oscillation now actually depend upon the quantum effects (albeit in an admittedly almost imperceptible manner). The solution (42) with k=1k=-1 describes an open universe emerging from an initial singularity at $t=2\sqrt{\delta}/c^{2}$, and (43) deals with a flat universe, which is also born from a singularity at $t=0$.
    (ii) $\delta=0$. This case as well the next one can only be realized in the presence of a quantum potential. The equation (33) reduces to ${\dot{x}}^{2}+kc^{2}=0$ which has real-valued solutions either when k=1k=-1 ($x(t)=\pm c(t-t_{0})$) or $k=0$ ($x={\rm const}$).
@@
    and the universe must be necessarily open (k=1k=-1).
@@ 5.2 Three is a company: $N=3$ and the Hamilton-Jacobi equation
Now, what kind of fields of matter should fill the third universe? In this subsection we will consider three cases: (I) there are no fields of matter ($\rho_{3}=p_{3}=0$) and the third universe is, like the second one, governed purely by a quantum potential; (II) a universe filled with matter whose parameter of state is w=1/3w=-1/3 (and so $\rho_{3}=\sigma^{2}/a_{3}^{2}$); (III) a hot universe filled by radiation with $w=1/3$ (here $\rho_{3}=R^{2}/a_{3}^{4}$). The reason we chose them specifically is because in these three cases the equations turns out to be completely integrable and leading to solutions with a strikingly similar behaviour. It is crucial to keep in mind that unlike the previously discussed cases, when $N>2$ the equations (16), (17) generally cease to be completely integrable, so the importance of those rare cases with known exact solutions cannot be overestimated. In fact, in the remainder of this section we will discuss both the solutions and the method of obtaining them.
@@
$\begin{array}[]{l}\displaystyle{S(t,r,\phi,\alpha,E)=-Et+\kappa_{2}\sqrt{2}\int d\phi\sqrt{\alpha-v(\phi)}+}\\ \\ \displaystyle{+\kappa_{1}\sqrt{2\alpha}\left(\sqrt{\frac{Er^{2}}{\alpha}-1}-\arctan\sqrt{\frac{Er^{2}}{\alpha}-1}\right),}\end{array}$ (54)
@@
The behaviour of the solution (56) explicitly depends upon the choice of $v(\phi)$ and will be different in cases (I), (II) and (III). On the other hand, we can still surmise asymptotic behaviour of the universes #2\#2 and #3\#3 without actually taking the integrals in (56). Indeed, using (47) we can calculate the Hubble parameter:
@@
and $v(\phi)$ is given by (50). Note, that except for the pathologically small values of constant $R\to 0$, the function $v(\phi)$ remains strictly positive in between the stop points $\phi_{1}$ and $\phi_{2}$ that satisfy the condition $0<\phi_{1}<\phi_{2}<\pi/4$, thus making sure the solutions (60) remain real valued.